# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

### v2

@@ 5.1 Data and Feature Construction
  • Rolling Volatility: A 20-day rolling standard deviation, $\sigma_{t}=\sqrt{\frac{1}{20}\sum_{k=1}^{20}(r_{t-k}-\bar{r})^{2}}$, to capture recent realized volatility.
@@ 5.2.1 Temporal Conformal Prediction (TCP)
Our proposed model first estimates the conditional quantiles, $\widehat{q}_{\tau}(t)$, using a gradient-boosted tree model, $\mathcal{F}_{\tau}$, trained on the feature set described above. The prediction interval is then formed by taking these quantile predictions and adjusting them with an adaptive conformal threshold, $C_{t}$. This threshold is updated at each time step based on the observed coverage error, as detailed in Section 4. The entire process is performed sequentially on a rolling basis with a window of $w=252$ days.
@@ 5.2.2 Benchmark Models
We compare TCP against three widely used models:
@@
    Quantile Regression (QR): A static version of the same gradient-boosted tree model used in TCP, but trained once on the entire dataset without any temporal adaptation or conformal calibration.
@@
    GARCH(1,1): A standard parametric model where the conditional variance $\sigma_{t}^{2}$ is modeled as $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\sigma_{t-1}^{2}$. We assume conditional normality to form prediction intervals.
@@
    Historical Simulation (Hist): A non-parametric approach where the prediction interval is formed by taking the empirical quantiles of the returns from a rolling 252-day window.
@@ 5.3 Evaluation Framework
5.4 Implementation Details
For transparency and reproducibility, we conclude our methodology section by providing details on the software, libraries, and training protocols used in our experiments.
All models were implemented in Python (v3.11) using the LightGBM library for the gradient-boosted tree models. The training protocol involves an initial warm-up period for each model, after which predictions are generated sequentially. For reproducibility, the complete code pipeline and data are made publicly available at github.com/agnivibes/tcp. The per-step computational complexity of TCP is approximately $O(w\log w)$ (Ke et al., 2017), driven by retraining the quantile regression model on a window of size $w$, making it highly suitable for real-time applications.
@@ 6.   Results and Discussion
Having detailed our methodology, we now present and discuss the results of our empirical investigation. This section analyzes the quantitative performance of all models, visualizes their behavior during a market crisis, and examines the robustness of our proposed TCP framework via hyperparameter sensitivity analysis.
Our empirical evaluation is designed to assess the models on two primary criteria: their ability to achieve the nominal 95% coverage target and the sharpness (width) of their prediction intervals. The main quantitative results are presented in Table 1, while Figure 1 provides a qualitative visualization of the models’ adaptive behavior during a period of extreme market stress.
@@
SP500 TCP 0.861 2.661 1449
QR 0.971 2.229 1701
GARCH 0.827 3.051 1670
Hist 0.931 5.058 1468
BTC-USD TCP 0.885 9.611 1449
QR 0.969 7.849 1701
GARCH 0.853 11.391 1670
Hist 0.944 18.058 1468
Gold TCP 0.881 2.202 1449
QR 0.969 1.800 1701
GARCH 0.837 2.618 1670
Hist 0.933 4.024 1468
Figure 1: A comparison of 95% prediction intervals from all four models for S&P 500 daily returns during the COVID-19 market crash (February-April 2020). The adaptive nature of TCP is clearly visible as its interval band widens in response to increased volatility in March and narrows as the market stabilizes in April.
@@ 6.1 Main Findings: Adaptiveness and Calibration
The central finding of our study is twofold. First, static machine learning models like Quantile Regression, while producing sharp intervals, are fundamentally unreliable for risk forecasting due to their poor calibration. As shown in Table 1, the QR model consistently overshoots the 95% target, achieving  97% coverage across all assets. In risk management, such systematic miscalibration is a critical flaw, as it leads to an inefficient allocation of capital.
Second, our proposed TCP framework demonstrates a clear and superior ability to adapt to changing market conditions. This is powerfully illustrated in Figure 1. During the low-volatility period of February 2020 for S & P 500, TCP’s interval is appropriately narrow. As the market crashes in March, the interval intelligently and dramatically expands to account for the heightened risk. Crucially, it begins to contract again in April as the market stabilizes. This dynamic behavior is precisely what is required of a robust risk management system. In contrast, the QR model’s interval remains static and inefficiently wide, while the GARCH and Historical Simulation models exhibit clumsy and less responsive adaptations. Similar visualizations for BTC-USD and Gold, which show consistent findings, are provided in Appendix B.
While TCP’s empirical coverage of 86-88% undershoots the 95% target in this experiment, its adaptive mechanism, driven by the modified Robbins-Monro scheme, represents a principled attempt at calibration that is entirely absent in the static QR model. This ability to navigate the coverage-sharpness tradeoff in real-time is the core contribution of our framework. In summary, the results reveal that TCP is a more reliable and trustworthy tool for real-world financial forecasting, where adaptability to non-stationarity and adherence to specified risk tolerances are paramount.
6.2 Hyperparameter Sensitivity
To evaluate the robustness of TCP, we performed a sensitivity analysis on its key hyperparameters: the window size $w$ and the initial learning rate $\gamma_{0}$. We tested TCP on the S&P 500 dataset with window sizes $w\in\{100,252,500\}$ and learning rates $\gamma_{0}\in\{0.005,0.01,0.05\}$. The results are summarized in Table 2. This analysis was repeated for BTC-USD and Gold, yielding similar conclusions about the model’s stability; these additional results are provided in Appendix C.
Table 2: TCP Sensitivity Analysis on S&P 500 Data
@@ 6.2 Hyperparameter Sensitivity
Window Size ($w$) Learning Rate ($\gamma_{0}$) Coverage Interval Width
100 0.005 0.8463 2.6521
100 0.010 0.8569 2.6288
100 0.050 0.8879 2.5847
252 0.005 0.8496 2.7095
252 0.010 0.8613 2.6611
252 0.050 0.8958 2.5923
500 0.005 0.8524 2.7712
500 0.010 0.8631 2.7309
500 0.050 0.8965 2.6614
The analysis shows that TCP’s performance is remarkably stable. While a larger window size ($w=500$) slightly increases interval width as it incorporates more historical data, the coverage remains consistent. Similarly, a higher learning rate ($\gamma_{0}=0.05$) allows for faster adaptation, pushing the coverage closer to 90% without causing instability. Across all settings, TCP maintains a strong balance between coverage and sharpness, confirming that its effectiveness is not an artifact of fine-tuning. This robustness is a critical feature for practical deployment, as it reduces the need for costly hyperparameter optimization.
@@ 7.   Conclusion and Future Work
In this paper, we introduced Temporal Conformal Prediction (TCP), a flexible and adaptive framework for generating prediction intervals in financial time series. Our method combines the strengths of quantile regression with conformal calibration, and crucially adapts over time using a decaying learning rate to correct for miscalibration. Unlike traditional models, TCP does not assume stationarity or rely on rigid parametric forms, making it well-suited to handle the complex, noisy, and often non-i.i.d. nature of financial returns. Unlike existing adaptive conformal approaches (Gibbs et al., 2021), TCP explicitly leverages temporal exchangeability and a modified Robbins-Monro scheme to handle the unique non-stationarity of financial time series. This makes it particularly suited for regulatory applications where robustness to market regime shifts is critical.
We benchmarked TCP against three widely used alternatives: standard Quantile Regression (QR), GARCH models, and Historical Simulation. Our results, across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold), highlight a crucial finding: static machine learning models like QR, while producing sharp intervals, are fundamentally miscalibrated and systematically overshoot the nominal 95% coverage target. In contrast, TCP’s adaptive mechanism demonstrates a clear and superior ability to adjust its intervals in response to market volatility, offering a more principled approach to achieving the desired coverage. Furthermore, a detailed sensitivity analysis confirmed that TCP’s performance is robust across a range of hyperparameter settings, underscoring the stability of the framework.
One of the key takeaways is that TCP offers a principled way to navigate the coverage–sharpness tradeoff, which is central to real-world forecasting applications. In practice, achieving properly calibrated intervals is paramount for interpretability, regulatory compliance, and capital efficiency. TCP’s distribution-free design aligns with Basel III/IV and SEC stress-testing requirements, which mandate robust risk assessments under varying market conditions. By providing valid intervals without parametric assumptions, TCP could streamline compliance workflows while improving capital allocation efficiency. By design, TCP fills a growing need for distribution-free, adaptive forecasting tools in financial modeling, a high-impact domain given increased scrutiny around risk assessments and regulatory robustness.
@@ 7.1 Future Work
Building on the foundation established in this work, we identify several avenues for future exploration. While our study focuses on univariate daily returns, several promising extensions remain. First, TCP can be generalized to multivariate or portfolio‐level forecasting by modeling joint or conditional quantiles This is where deep copula models (Aich et al., 2025) could be used to capture complex, non-linear dependencies between assets. Second, richer quantile estimators, such as neural networks or advanced tree‐based methods using additional market features, may further improve interval sharpness and adaptability. Third, it would be valuable to benchmark TCP against volatility models that leverage realized‐volatility inputs, for example true Realized-GARCH fitted on intraday realized variances or Realized-GARCH variants using rolling-window proxies. Fourth, a particularly promising avenue is applying TCP to the detection of financial bubbles; the behavior of the interval bounds themselves could serve as a novel, real-time indicator of market exuberance, while the model’s standard crash signal could be used to forecast the subsequent implosion, providing a new perspective on the work of Phillips et al. from 2018. Furthermore, the framework can be adapted from general risk monitoring to event-specific forecasting, such as predicting the likelihood of extreme monthly drawdowns (e.g., $>$10% market declines) by aggregating the model’s daily signals. Finally, deploying TCP in live trading simulations or real-time risk‐management systems will help assess its operational robustness, latency, and scalability. We leave these directions for future exploration.
Finally, the framework’s principles are highly applicable beyond finance, wherever calibrated uncertainty quantification is needed for non-stationary processes. In marketing, for instance, TCP could reframe forecasting as a risk management problem, providing robust prediction intervals for demand forecasting and inventory management, or quantifying the uncertain returns from marketing mix models and budget allocations. In biostatistics, its ability to handle non-stationarity is ideal for real-time patient monitoring in ICUs, forecasting blood-glucose ranges for chronic disease management, and adapting personalized drug dosing. Further applications exist in renewable energy for forecasting power output to ensure grid stability and in sports analytics for dynamic, in-game strategic decision-making and injury risk assessment. Exploring these domains would underscore TCP’s potential as a foundational tool for adaptive forecasting.

### v3

@@ 5.1 Data and Feature Construction
  • Rolling Volatility: $\sigma_{t}=\sqrt{\frac{1}{20-1}\sum_{k=1}^{20}\big(r_{t-k}-\bar{r}_{t,20}\big)^{2}}\quad\text{(rolling sample std, ddof=1)}$ where $\bar{r}_{t,20}=\frac{1}{20}\sum_{k=1}^{20}r_{t-k}$.
@@
In implementation we use the rolling sample standard deviation via pandas (.rolling(20).std, ddof=1=1).
@@ 5.2.1 Temporal Conformal Prediction (TCP)
Our proposed model first estimates the conditional quantiles, $\widehat{q}_{\tau}(t)$, using a gradient-boosted tree model, $\mathcal{F}_{\tau}$, trained on the feature set described above. The prediction interval is formed from the forecaster’s two quantiles and the split-conformal threshold $C_{t}$ computed on a disjoint calibration slice (Alg. 1). In the TCP-RM variant, $C_{t}$ is further updated online via (1). This threshold is updated at each time step based on the observed coverage error, as detailed in Section 4. The entire process is performed sequentially on a rolling basis with a window of $w=252$ days.
@@ 5.2.2 Benchmark Models
    Quantile Regression (QR): Same gradient-boosted quantile forecaster as TCP, evaluated sequentially in a rolling, out-of-sample manner on a trailing window of $w_{\mathrm{tr}}{=}192$ observations (no conformal calibration).
@@
    Fixed-parameter GARCH(1,1) (EWMA): Conditional variance updated recursively
    $\sigma_{t}^{2}=\omega+\alpha r_{t-1}^{2}+\beta\,\sigma_{t-1}^{2},\qquad(\omega,\alpha,\beta)=(10^{-6},\,0.05,\,0.9),$
    with coefficients held fixed and updated sequentially using only past data (no rolling re-estimation). One-step-ahead $(1-\alpha)$ intervals assume conditional normality with zero mean, i.e., $z_{1-\alpha/2}\sigma_{t},\;z_{1-\alpha/2}\sigma_{t}\,]$.
@@
    Historical Simulation (Hist): Non-parametric empirical $(\alpha/2,\,1-\alpha/2)$ quantiles from a rolling 252-day window.
@@ 6.  RESULTS AND DISCUSSION
We now assess Temporal Conformal Prediction (TCP) and our main contribution TCP-RM (TCP with a Robbins–Monro online calibration layer) against strong baselines (QR, GARCH, Historical Simulation) on three asset classes: S&P 500, BTC-USD, and Gold. The goal is to test whether the online calibration delivers near-nominal coverage under distribution shift while preserving sharpness.
Evaluation centers on two criteria: (i) empirical coverage versus the nominal target $1-\alpha=0.95$, and (ii) average interval width as a measure of sharpness. Table 1 summarizes full-sample performance; Figure 1 visualizes behavior during the COVID-19 crash. We also report a hyperparameter sensitivity study (window size $w$, stepsize $\gamma_{0}$) to assess robustness of TCP-RM.
@@
SP500 TCP 0.9441 3.5975 1449
TCP-RM 0.9469 3.6068 1449
QR 0.8628 2.686 1509
GARCH 0.8269 3.0505 1670
Hist 0.9312 5.0575 1468
BTC-USD TCP 0.9586 11.2346 1449
TCP-RM 0.9593 11.2374 1449
QR 0.8960 9.8446 1509
GARCH 0.8533 11.3908 1670
Hist 0.9441 18.0576 1468
Gold TCP 0.9427 2.6630 1449
TCP-RM 0.9448 2.6696 1449
QR 0.8761 2.2394 1509
GARCH 0.8365 2.6185 1670
Hist 0.9326 4.0238 1468
Figure 1: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for S&P 500 daily returns during the COVID-19 market crash (Feb–Apr 2020). Shaded bands show the interval; the red dots mark days where the realized return falls outside the 95% interval (miscoverage). TCP/TCP-RM bands widen rapidly into the March spike and contract in April, illustrating adaptiveness. Analogous crisis-window panels for BTC-USD and Gold appear in Appendix B.
@@ 6.1 Main Findings: Adaptiveness and Calibration
Three findings emerge clearly from Table 1 and the crisis-window visualization (Fig. 1):
  1. 1.
    (i) Calibration vs. sharpness trade-off. The rolling ML baseline (QR) produces the narrowest intervals across all assets but is materially under-calibrated relative to the 95% target: SP500 86.3%, BTC 89.6%, Gold 87.6%. This under-coverage is the price for sharpness (e.g., SP500 average width 2.686 versus TCP’s 3.598). In other words, rolling QR concentrates mass too aggressively in volatile regimes.
  2. 2.
    (ii) TCP/TCP-RM achieve near-nominal coverage with substantially tighter bands than Historical Simulation. TCP coverage is close to the target on all assets (SP500 94.4%, BTC 95.9%, Gold 94.3%), while keeping intervals 30–38% narrower than Historical Simulation: SP500: 3.598 vs 5.058 (-28.9%); BTC: 11.235 vs 18.058 (-37.8%); Gold: 2.663 vs 4.024 (-33.8%). Thus, TCP delivers a better calibration–sharpness balance than Hist, which attains decent coverage but with overly conservative widths.
  3. 3.
    (iii) Robbins–Monro improves calibration at essentially no cost in width. TCP-RM yields a small positive lift in coverage relative to TCP with negligible width change on all assets (e.g., SP500: +0.28 percentage points coverage, +0.009 width). This supports the utility of the online offset for fine-tuning calibration without sacrificing sharpness.
GARCH under-covers consistently ($\approx$0.83–0.85) while not being as sharp as QR, suggesting its conditional-Gaussian volatility path is too slow to adapt to return asymmetries and tail thickness during stress.
Crisis-window behavior. Figure 1 (SP500, Feb–Apr 2020) shows how each method behaves during the shock. The red dots mark days where the realized return falls outside the reported 95% prediction interval (i.e., miscoverage events). TCP bands widen abruptly in early March, remain elevated through the peak volatility in mid–late March, and begin to contract in April as conditions stabilize. TCP and TCP-RM deliver similar behavior and, in this short 55-day window, both achieve coverage of $\approx 0.80$ with average widths $\approx 11.55$–11.57. Rolling QR widens only modestly, keeps much tighter bands (width =7.69=7.69), and shows the most red dots around the extremes, consistent with its lower window coverage (0.709). GARCH ramps up more smoothly, reflecting its volatility recursion, and lands at the same window coverage as TCP (0.800) with narrower bands (width =9.66=9.66). Historical Simulation reacts more slowly and is the most under-covered here (0.655), despite having the narrowest bands in this window (width =5.91=5.91); this lag is expected from its long empirical lookback. Overall, the crisis snapshot illustrates a general pattern under abrupt regime change: all methods temporarily under-cover, but TCP/TCP-RM adapt fastest by inflating bands right when volatility jumps, while QR remains sharp at the cost of additional misses. This visual evidence complements the full out-of-sample results in Table 1, where TCP/TCP-RM recover to near-nominal calibration over the full evaluation horizon. Analogous crisis-window panels for BTC-USD and Gold are reported in the Appendix B.
Implementation note (comparability). All learning-based methods use the same feature set. TCP/TCP-RM are evaluated out-of-sample with a rolling window of 252 points split into 192 training and 60 calibration points. Rolling QR is also evaluated out-of-sample using a 192-point trailing training window. The slight differences in the “Predictions” counts (e.g., 1509 for QR vs. 1449 for TCP/TCP-RM) reflect different warm-up requirements (QR does not reserve a calibration slice). Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: QR is sharp but under-calibrated; TCP/TCP-RM are near-nominal and materially sharper than Historical Simulation.
6.2 Hyperparameter Sensitivity (SP500, TCP-RM)
We stress–test the adaptive layer by varying the rolling window $w$ and the initial step size $\gamma_{0}$ for TCP-RM(Table 2). For comparability we fix the calibration slice at $|\mathcal{C}|=40$ and sweep $w\in\{100,252,500\}$ and $\gamma_{0}\in\{0.005,0.01,0.05\}$ on SP500. Results for BTC-USD and Gold exhibit the same pattern and are reported in Appendix D.
Table 2: TCP-RM sensitivity on S&P 500 with $|\mathcal{C}|=40$ (target 95%).
@@ 6.2 Hyperparameter Sensitivity (SP500, TCP-RM)
Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
100 0.005 0.9557 4.5188
100 0.010 0.9557 4.5202
100 0.050 0.9550 4.5315
252 0.005 0.9469 3.4605
252 0.010 0.9469 3.4673
252 0.050 0.9524 3.4878
500 0.005 0.9559 3.2627
500 0.010 0.9659 3.2815
500 0.050 0.9759 3.3970
What we learn. (i) Increasing $w$ consistently sharpens intervals: widths fall from $\approx 4.52$ (w=100w{=}100) to 3.263.263.403.40 (w=500w{=}500), while coverage stays at or above the target band. (ii) At a fixed $w$, larger $\gamma_{0}$ nudges coverage upward with a small width trade-off; e.g., at w=252w{=}252, $\gamma_{0}{=}0.050$ lifts coverage from $0.9469$ to $0.9524$ (+0.55pp) with a $\sim 0.03$ increase in width; at w=500w{=}500, very large $\gamma_{0}$ overshoots the target (0.9759).
Takeaway. TCP-RM is robust to reasonable choices of $(w,\gamma_{0})$. For the main experiments, we retain w=252w{=}252 and $\gamma_{0}{=}0.01$, which stay near the nominal level while keeping intervals tight; if one prefers even sharper intervals, w=500w{=}500 with $\gamma_{0}\in[0.005,0.01]$ is a good choice without material over-coverage.
We also perform sensitivity analysis on TCP. The results are given in Appendix E.
6.3 Visualization for TCP vs. TCP-RM on S&P 500 (95% intervals)
In this section, we present the TCP vs TCP-RM trace plot for S&P 500.
Figure 2: TCP vs. TCP-RM on S&P 500 (95% intervals). Top: daily returns with two‐sided prediction bands for TCP (blue) and TCP-RM (purple). Middle: 30-day rolling empirical coverage; the dashed line marks the 95% target. Bottom: evolution of conformal thresholds: split‐conformal level $C_{\text{split}}$, online Robbins–Monro offset $C_{\text{RM}}$, and the effective total $C_{\text{eff}}=C_{\text{split}}+C_{\text{RM}}$. During the COVID-19 crash the bands widen sharply, the rolling coverage briefly departs from the 95% target and then re-enters the band, and $C_{\text{eff}}$ spikes before decaying as conditions stabilize.
In Fig. 2, TCP and TCP-RM behave similarly in tranquil periods, but during stress the effective threshold $C_{\text{eff}}$ (bottom panel) rises quickly, producing wider bands (top) and steering the 30-day rolling coverage back toward the 95% target (middle). The Robbins–Monro offset $C_{\text{RM}}$ remains near zero except around regime shifts, where it provides the extra adjustment that closes the gap to target coverage. This visual evidence complements Table 1: on S&P 500, TCP-RM attains $\approx$95% coverage with only a marginal width increase relative to TCP
@@ 7.  CONCLUSION AND FUTURE WORK
This paper introduced Temporal Conformal Prediction (TCP), a simple and effective framework for calibrated prediction intervals in nonstationary time series. TCP combines a rolling, exchangeability‐aware split–conformal layer with a modern quantile forecaster, and the TCP–RM variant adds a single Robbins–Monro offset that adjusts the conformal threshold online using observed coverage errors.
Across three asset classes, TCP/TCP–RM deliver a strong calibration–sharpness balance: coverage is close to the 95% target while intervals are materially narrower than Historical Simulation. In contrast, the rolling QR baseline attains the narrowest bands but undershoots the target coverage, especially in turbulent regimes. Crisis‐window visualizations around the COVID‐19 crash illustrate TCP’s intended adaptiveness: intervals widen rapidly as volatility spikes and contract as conditions stabilize, with red dots indicating miscoverage events. A targeted sensitivity analysis shows robustness to reasonable choices of rolling window and Robbins–Monro step size, and suggests default settings that are stable in practice.
Limitations. Like any sequential method, TCP can temporarily under‐cover during abrupt regime shifts before its calibration catches up. Performance also inherits the quality of the base quantile forecaster and the choice of windowing; extremely heavy tails or structural breaks can stress any fixed feature set. Finally, while TCP–RM improves calibration at negligible width cost on average, very aggressive step sizes can overcorrect.
Practical implications. By providing distribution‐free intervals that adapt in time, TCP aligns naturally with backtesting and governance needs in risk management (coverage targets, exception tracking, and scenario robustness). The framework is lightweight (single threshold) and model‐agnostic (plugs into any quantile forecaster), making deployment straightforward in production pipelines.
@@ 7.1 Future Work
Several extensions follow naturally:
  1. 1.
    Multivariate and portfolio TCP. Generalize to joint/conditional quantiles for baskets of assets; couple TCP with copula or factor models to control portfolio‐level coverage.
  2. 2.
    Conditional and localized coverage. Target coverage conditional on covariates (e.g., volatility state) and explore adaptive nonconformity scores that emphasize tail risk or signed errors.
  3. 3.
    Richer forecasters and features. Swap the base learner for neural quantile models or boosted variants with realized‐volatility inputs (intraday RV), order‐flow, or macro features; assess benefits for sharpness at fixed coverage.
  4. 4.
    Change–point aware adaptation. Trigger temporary step–size boosts or threshold projections when a regime break is detected to shorten the transient under/over‐shoot in coverage.
  5. 5.
    Heavy‐tail robustness. Investigate EVT‐inspired nonconformity scores or asymmetric penalties to better stabilize performance under extreme shocks.
  6. 6.
    Decision integration. Map TCP intervals to VaR/ES, margin, and capital rules; evaluate utility/capital efficiency vs. conservative baselines in realistic backtests.
  7. 7.
    Beyond finance. TCP’s adaptiveness is broadly applicable wherever nonstationarity and calibrated uncertainty matter:
    • Biostatistics/medicine: ICU vitals monitoring and glucose‐range forecasting with patient–specific adaptation; adaptive dosing intervals under drift.
    • Epidemiology: Calibrated intervals for nowcasts/forecasts of incidence and $R_{t}$ that widen under reporting delays or testing-policy shifts, plus adaptive risk bands for hospitalization/ICU load as transmission regimes change.
    • Genetics/Genomics: Uncertainty bands for polygenic risk scores and eQTL effect sizes that remain calibrated under population/covariate shift; interval forecasts for gene expression (bulk or single-cell) with batch effects and sequencing-depth drift; calibrated pathogenicity scores for rare variants.
    • Sports analytics: Game-state–aware intervals for win probability, expected goals/points, and player performance that adapt to tempo and lineup changes; injury-risk ranges for load management with rapid recalibration after shocks (e.g., in-game injuries or schedule congestion).
    • Renewable energy: Wind/solar generation intervals that expand/contract with weather regimes to aid grid balancing.
    • Operations/marketing: Demand and lead‐time intervals for inventory control; calibrated lift ranges in marketing‐mix models under campaign drift.
    • Transportation and climate: Travel‐time intervals under traffic regime shifts; local hazard/risk bands for extreme weather.
In sum, TCP offers a compact, distribution‐free, and adaptive route to calibrated uncertainty quantification in the presence of temporal drift. Its empirical performance, visual adaptiveness in crisis windows, and robustness to sensible hyperparameters suggest it is a practical default for risk forecasting and a promising building block for broader sequential decision systems.
8.  DATA AND CODE AVAILABILITY
The financial returns dataset and all code used in this paper are publicly available at: https://github.com/agnivibes/temporal-conformal-prediction-tcp.